# A hybrid quantum solver for the Lorenz system

### v2

@@ 1 Introduction
Dynamical systems theory, along with its numerical and simulated counterparts, has been employed to study various phenomena such as weather patterns, population dynamics, economic trends, the flow of chemical and biological elements, etc Thompson (2002). Whether it is the macroscopic dynamics of a temperature-pressure system or the microscopic complexities of a nonlinear system with bifurcations and irreversible physics, mathematical modelling, based on dynamical system theory, offers a valuable tool for analysis.
@@
When a single iteration of the Euler method is solved using a quantum algorithm, there are two intrinsic uncertainties. The initial state can only be prepared with some certainty, and the solution computed by a quantum algorithm is really a random sample (given by the probabilities). Hence, there is no reason to believe that the trajectory of the classical solution will be the same as the trajectory of the solution computed using a quantum method, even if the two start points are the same.
Suppose we replace the equation-solving step in the Euler method with a quantum algorithm; given the new source of uncertainty, the following questions arise: i) does the chaos in the system increase when a quantum subroutine is used in each iteration? ii) does the system still have the butterfly attractor (for the choice of parameters given above)? We answer these questions in the affirmative under the assumption that it is possible to prepare the state exactly and recover the solution exactly. These assumptions can only be justified with sufficient advances in Quantum random-access memory (QRAM) and quantum tomography techniques Lanyon (2017); Cramer (2010).
One of the earliest attempts, to solve nonlinear differential equations using a quantum approach, is due to Leyton (2008). They use the forward Euler method and multiple copies of an initial state, which are evolved according to the linear Euler system. This algorithm scaled poorly as a function of the time step, as shown in Berry Berry (2014), which goes on to show that a far more efficient quantum procedure can be obtained for linear differential equations. For this, they compute the bounds on the condition number of the matrix. Liu (2021) gives an efficient quantum algorithm for dissipative nonlinear differential equations.
@@ 2.2 Inclusion of nonlinear terms
This inversion can be done classically or using HHL Lloyd (2020) as in Tennie (2023). HHL assumes Hermitian matrix, therefore we solve for ${\mathcal{A}}_{NL}{\mathcal{W}}_{NL}=B_{NL}$ where ${\mathcal{A}}_{NL}=\begin{bmatrix}A_{NL}&\mathcal{O}_{8}\\ \mathcal{O}_{8}&A_{NL}^{\dagger}\end{bmatrix}$, ${\mathcal{W}}_{NL}=(W,W)$ and $B_{NL}=(b_{NL},b_{NL})$ and with $\mathcal{O}_{8}$ a $8\times 8$ matrix with all 00 entries.
We need $4$ qubits to represent the input $b_{NL}$, one qubit for controlled rotation, and say 4 qubits for the clock register (needed for simulation of the Hamiltonian). So, a total of 9 qubits are needed if HHL is used to solve $A_{NL}W=b_{NL}$. The precision increases with the increase in the number of clock qubits. The condition number for $\mathcal{A}_{NL}$ is 3.03 for $h$, step size, of $0.01$.
@@
On the other hand, Berry (2014) has an efficient quantum procedure for linear differential equations, while in the present work, we consider a nonlinear system, so the approach here is not directly comparable with the results there.
@@
To solve (22), we begin from an initial point and use iterative methods to find the optimal solution. While numerous classical methods are available Butcher (1996), we will use a variational quantum algorithm called VQLS.
@@ 3.3 Ansatz
An Ansatz is a quantum circuit used in variational algorithms. The Ansatz acts as an initial guess for the quantum state and is usually defined by a set of parameters. These parameters are then adjusted to minimize a cost function, allowing the Ansatz to approximate the intended quantum state or solve a specific computational problem. The flexibility and adaptability of the Ansatz are essential for the success of variational quantum algorithms.
The Strongly Entangling Layers (SEL) Ansatz is a type of parameterized quantum circuit used in variational quantum algorithms, such as the VQE and Quantum Approximate Optimization Algorithm (QAOA). This Ansatz is designed to introduce a high degree of entanglement between qubits while maintaining a relatively simple and regular structure, making it a popular choice for many quantum applications. Let’s describe the structure of the SEL Ansatz.

### v3

@@ 1 Introduction
Dynamical systems theory, along with its numerical and simulated counterparts, has been employed to study various phenomena such as weather patterns, population dynamics, economic trends, the flow of chemical and biological elements, etc Thompson (2002). Whether it is the macroscopic dynamics of a temperature-pressure system or the microscopic complexities of a nonlinear system with bifurcations and irreversible physics, mathematical modeling, based on dynamical system theory, offers a valuable tool for analysis.
@@
When solving a single iteration of the Euler method using a quantum algorithm, there are two sources of uncertainty. The initial state can only be prepared with a certain level of precision, and the solution computed by a quantum algorithm is inherently probabilistic, resulting in a random sample rather than a deterministic value. However, it is important to note that this randomness does not introduce new features or chaotic behavior into the system. The differences in trajectories between classical and quantum solutions arise from the inherent probabilistic nature of quantum algorithms, not from any additional characteristics of the Lorenz system itself. Both methods—classical and quantum—can provide approximate solutions, but the quantum method may do so more efficiently by exploiting quantum randomness, similar to how stochastic classical methods use randomness.
Suppose we replace the equation-solving step in the Euler method with a quantum algorithm; given the new source of uncertainty, the following questions arise: i) does the chaos in the system increase when a quantum subroutine is used in each iteration? ii) does the system still have the butterfly attractor (for the choice of parameters given above)? We answer these questions in the affirmative under the assumption that it is possible to prepare the state exactly and recover the solution exactly. These assumptions can only be justified with sufficient advances in Quantum random-access memory (QRAM) and quantum tomography techniques Lanyon (2017); Cramer (2010)
One of the earliest attempts, to solve nonlinear differential equations using a quantum approach, is due to Leyton (2008). They use the forward Euler method and multiple copies of an initial state, which are evolved according to the linear Euler system. This algorithm scaled poorly as a function of the time step, as shown in Berry (2014), which goes on to show that a far more efficient quantum procedure can be obtained for linear differential equations. For this, they compute the bounds on the condition number of the matrix. Liu (2021) gives an efficient quantum algorithm for dissipative nonlinear differential equations.
@@ 2.2 Inclusion of nonlinear terms
This inversion can be done classically or using HHL Lloyd (2020) as in Tennie (2023). HHL assumes Hermitian matrix, therefore one has to solve for ${\mathcal{A}}_{NL}{\mathcal{W}}_{NL}=B_{NL}$ where ${\mathcal{A}}_{NL}=\begin{bmatrix}\mathcal{O}_{8}&A_{NL}\\ A_{NL}^{\dagger}&\mathcal{O}_{8}\end{bmatrix}$, ${\mathcal{W}}_{NL}=(\mathbf{O}_{8},W)$ and $B_{NL}=(b_{NL},\mathbf{O}_{8})$ and with $\mathcal{O}_{8}$ a $8\times 8$ matrix with all 00 entries, and $\mathbf{O}_{8}$ is an 8-dimensional column vector of zeros. We need $4$ qubits to represent the input $b_{NL}$, one qubit for controlled rotation, and say 4 qubits for the clock register (needed for simulation of the Hamiltonian). So, a total of 9 qubits are needed if HHL is used to solve $A_{NL}W=b_{NL}$. The precision increases with the increase in the number of clock qubits. The condition number for $\mathcal{A}_{NL}$ is 3.03 for $h$, step size, of $0.01$.
Although the HHL and VQLS methods are both prominent algorithms for solving linear systems, they present distinct advantages and limitations. The HHL algorithm is known for its theoretical exponential speedup over classical methods for certain classes of linear systems, particularly those matrices $A$ with the sparse and well-conditioned Harrow (2009); Childs (2017). However, HHL is highly susceptible to quantum gate errors and qubit decoherence, which significantly impair its performance on current quantum hardware Campbell (2017); Preskill (2018). Additionally the simulation of HHL is computationally intensive.
In contrast, the VQLS method employs simpler and often shorter quantum circuits, making it more feasible on available noisy intermediate-scale quantum (NISQ) hardware Bravo-Prieto (2023); Cerezo (2021). While VQLS may not achieve the same exponential speedup as HHL, its variational approach facilitates noise mitigation, resulting in greater robustness and reliability in noisy quantum environments Endo (2018); McArdle (2020). Additionally, VQLS requires fewer quantum resources by combining quantum subroutines with classical optimization, making it particularly suited to NISQ devices, where quantum coherence time is limited McClean (2017); Benedetti (2019).
@@
On the other hand, Berry (2014) has an efficient quantum procedure for linear differential equation, while in the present work, we consider a nonlinear system, so the approach here is not directly comparable with the results there.
@@
To show the efficiency of the proposed method, we compare the results of our proposed method and the Euler method, which is a non-linear approach, to solve the Lorenz system. We use the same number of iterations, step size, and initial point. We plotted this compression in Figure 3.
Figure 3: The starting point is $(1,1,1)$.
Encouragingly, the solutions obtained from our proposed linear system approach are similar to the those from nonlinear Euler method. This result demonstrates that although our method is based on solving a linear system, it effectively captures the dynamics of the nonlinear Lorenz system. To solve (22), we begin from an initial point and use iterative methods to find the optimal solution. While numerous classical methods are available Butcher (1996), we will use a variational quantum algorithm called VQLS.
@@ 3.3 Ansatz
In variational algorithms, an Ansatz refers to an assumed initial form for the quantum state. This Ansatz is typically represented by a parameterized quantum circuit, which is used to prepare a trial state that can be optimized. The Strongly Entangling Layers (SEL) Ansatz is a type of parameterized quantum circuit used in variational quantum algorithms, such as the VQE and Quantum Approximate Optimization Algorithm (QAOA). This Ansatz is designed to introduce a high degree of entanglement between qubits while maintaining a relatively simple and regular structure, making it a popular choice for many quantum applications. Let’s describe the structure of the SEL Ansatz.
@@ 5.5 Error Analysis
5.6 Error for VQLS Method
Here, we briefly review some key concepts related to errors in the VQLS method Bravo-Prieto (2023). For the VQLS algorithm, the deviation between observable expectation values for the approximate solution $|x(\alpha_{\text{opt}})\rangle$ and the true solution $|x_{0}\rangle$ can be upper bounded based on the value of the cost function. Specifically, the error tolerance $\epsilon$ can be set before running the algorithm, where $\epsilon$ is defined as the trace distance between the exact and approximate solutions:
$\epsilon=\frac{1}{2}\text{Tr}\left(\left\||x_{0}\rangle\langle x_{0}|-|x(\alpha_{\text{opt}})\rangle\langle x(\alpha_{\text{opt}})|\right\|\right)$ (35)
Moreover, it has been proven that the cost function satisfies the relation Bravo-Prieto (2023):
$C_{G}\geq\frac{\epsilon^{2}}{\kappa^{2}}$ (36)
where $\kappa$ is the condition number of the input matrix, and $C_{G}$ is the cost function defined by (27). By using the above formulas, we can observe that the difference between the exact solution and the quantum solution is related to $\kappa^{2}$ multiplied by the value of the objective function. This shows that the error between the quantum solution and the exact solution is inherently connected to the condition number of the matrix and the chosen cost function, allowing us to control and quantify the error before running the algorithm.
Several noise mitigation strategies are particularly well-suited to enhance the robustness of VQLS on NISQ devices due to the method’s variational structure and we can apply these methods. Techniques such as Zero-Noise Extrapolation (ZNE) and Measurement Error Mitigation (MEM) allow the VQLS to reduce the impact of noise without increasing the quantum circuit depth, making them practical for the shallow circuits typically used in variational algorithms Temme (2017); Kandala (2019). Additionally, the Variational Error Suppression (VES) technique can adaptively adjust the parameters within VQLS to minimize noise effects, enhancing solution accuracy even on noisy devices McClean (2017); Cerezo (2021). Together, these approaches demonstrate VQLS’s practical suitability for today’s quantum hardware while mitigating the effects of common noise sources.
@@ 5.8 Limitations
It is worth noting that the VQLS algorithm proposed for solving the Lorenz system aligns with existing implementations of quantum linear solvers for near-term quantum devices Bravo-Prieto (2023). In that study, a 10-qubit implementation of VQLS was performed on Rigetti’s Aspen-4 quantum computer to solve a Quantum Linear System Problem (QLSP) with a 1024 × 1024 matrix. This experiment utilized a specifically tailored ansatz incorporating $R_{y}(\alpha_{i})$ gates and computed the cost function by expanding the Hamiltonian in terms of Pauli operators to meet hardware constraints. The findings indicated that the cost function was minimized effectively on quantum hardware, closely matching simulated results and verifying that the solution to the linear system was obtained. This research, supports the feasibility of implementing our proposed method on current quantum devices.
5.9 Matrix Formulation for High-Dimensional Systems
To extend the proposed method to solve larger nonlinear systems of differential equations, we first need to ensure that the matrix $A_{NL}$ defined by (19) is compatible with quantum algorithms, which often require matrix dimensions to be powers of 2. If the original matrix $A_{NL}$ does not meet this requirement, we can pad it by adding an identity matrix or zero elements to reach the nearest $2^{n}\times 2^{n}$ dimension. This allows the system to be compatible with VQLS and enables efficient processing on quantum hardware.
Consider a general nonlinear system of ordinary differential equations given by:
$\frac{d\mathbf{u}}{dt}=\mathbf{f}(\mathbf{u}(t)),$ (37)
where $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),\ldots,u_{N}(t)]^{\top}$ is an $N$-dimensional vector of state variables, and $\mathbf{f}(\mathbf{u})=[f_{1}(\mathbf{u}),f_{2}(\mathbf{u}),\ldots,f_{N}(\mathbf{u})]^{\top}$ represents the nonlinear terms.
Using a timestep $h$, the discretized form at each time step $n$ can be written as:
$\mathbf{u}_{n+1}=\mathbf{u}_{n}+h\mathbf{f}(\mathbf{u}_{n}),$ (38)
which we rearrange as:
$\mathbf{u}_{n+1}-\mathbf{u}_{n}-h\mathbf{f}(\mathbf{u}_{n})=0.$ (39)
Now, we can rewrite the (39) into matrix form as:
$A_{NL}W=b_{NL},$ (40)
where: $A_{NL}$ is an $(N+M)\times(N+M)$ matrix (or nearest $2^{n}\times 2^{n}$ size if padding is applied), $W$ is an $(N+M)$dimensional vector containing state variables at the current and next time steps, and $b_{NL}$ denotes an $(N+M)$dimensional vector with known values. Now, we can apply the VQLS to solve the linear system given by (40).
@@ 6 Conclusion
VES Variational Error Suppression
ZNE Zero-Noise Extrapolation
MEM Measurement Error Mitigation
QLSP Quantum Linear System Problem